# Sparse Repellency for Shielded Generation in Text-to-Image Diffusion Models

### v2

@@ 4.6 Image Protection Benchmark
Figure 7: The images generated by EDMv2 in the first row are too close to existing images in the ImageNet-1k train set, which EDMv2 was trained on. SPELL adapts their diffusion trajectories to ensure that they maintain a protection radius. The images in third row, generated from the same seeds but with SPELL, are sufficiently different from their nearest neighbor.
@@ 5 Discussion
This work introduces sparse repellency (SPELL), a training-free post-hoc method to guide diffusion models away from a certain set of images. This both prevents repeating images that were already generated, thereby increasing diversity, and allows protecting a certain set of reference images, with applications like machine unlearning. SPELL can be applied to any diffusion model, whether it is class-to-image or text-to-image, and whether it is unconditional or classifier(-free) guided. We see two main ways in how future work can extend SPELL. First, SPELL can currently be proven to guarantee to generate images outside the shields if all shields are disjoint. If there are overlapping shields, there can be cases where the guarantee is not strict anymore, and the algorithm could be improved to find a direction that points out of the convex hull of all points, at the cost of decreased scalability. Second, we currently apply SPELL with respect to the $L_{2}$ distance inside the diffusion model’s latent spaces, which lead to visually different outputs. Using a distance inside a semantically structured space could lead to generating more semantically different images. This has the potential to further increase metrics like FID and recall, which evaluate generated images in semantic spaces.
@@ Appendix C Conservative Field Interpretation
The function $h(x)=\operatorname{ReLU}(\frac{r}{\|x\|}-1)x$ is the conservative field associated to the (family of) potential $H:\mathbb{R}^{d}\rightarrow\mathbb{R}$:
@@
Figure 8: Potential function whose gradient field is $\operatorname{ReLU}(\frac{r}{\|x\|}-1)x$, displayed for $x\in\mathbb{R}^{2}$. Repellence force is dynamic: closer to the center (i.e., when a diffusion trajectory is expected to be close to a protected image) it applies stronger gradients, as shown by the arrows, while outside the repellecy radius, it applies no gradient at all, letting the diffusion trajectory continue without any intervention.
@@ Appendix D Implementation Details and Hyperparameters
1 Input: Batch of latents $\{x_{t}^{(b)}\}_{b=1,\dots,B}$, set of shielded images $\{z_{k}\}_{k=1,\dots,K}$, radius $r$, $\lambda$
2 for $b=1,\dots,B$ do
@@
        $\triangleright$ Expected diffusion output without repellency
3
@@
5for $b=1,\dots,B$ do
6       $\vec{\Delta}_{b}\leftarrow 0$
7       for $k=1,\dots,K$ do $\triangleright$ Repel from the shielded set
@@ Appendix E Construction of the Soft-label CC12M Dataset
Appendix F Ablation: Repellency Strength Throughout the Generation
@@
Appendix G Image Protection on Large Datasets
@@
Appendix H Examples of Images Generated with Repellency
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ Appendix H Examples of Images Generated with Repellency
Appendix I Ablation: Changing the Guidance Weight

### v3

@@ 5.6 Image Protection at Scale
Figure 7: The images generated by EDMv2 in the first row are too close to existing images in the ImageNet-1k train set in the second row, which EDMv2 was trained on. SPELL ensures that they maintain a protection radius. The images in third row, generated from the same seeds but with SPELL, are sufficiently different from the ImagetNet neighbors in the second row, and also from their own nearest neighbors in the forth row.
@@ 6 Discussion
We introduce sparse repellency (SPELL), a training-free post-hoc method to guide diffusion models away from a set of images. SPELL increases diversity by preventing repeat generation, and allows protecting a set of given reference images. SPELL can be applied to any diffusion model, whether it is class-to-image or text-to-image, and whether it is unconditional or classifier(-free) guided, even at high resolution and arbitrarily sized reference sets.
We find three limitations of SPELL: First, theoretically, SPELL currently only guarantees generating images outside the shields if all shields are disjoint. If there are overlapping shields and the diffusion trajectory heads exactly into the middle of them, their repellency terms cancel. While we do not see this problem in practice even at the scale of 1.2M images, in theory, the algorithm could be improved: we can merge overlapping shields, or search for a direction that points out of the convex hull of all shields rather than their naive sum of parts, at the cost of scalability. Second, we currently apply SPELL with respect to the $L_{2}$ distance inside the diffusion models’ (VAE encoder) latent spaces, which lead to visually different outputs. Checking the distances inside a downstream semantically structured embedding space and propagating their directions back could lead to generating more semantically different images. Third, SPELL is applied to expected $\mathbb{E}[\mathbf{X}_{0}|x_{t}]$ under the forward joint distribution and not samples from $p_{0|t}$ required for Doob h transform. see Appendix A. This expectation will only correspond to the generative distribution if using the time-reversal diffusion, and not for example when using the probability flow ODE sampler.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix C Conservative Field Interpretation
The function $h(x)=\operatorname{\tiny\sigma_{\tiny relu}}(\frac{r}{\|x\|}-1)x$ is the conservative field associated to the (family of) potential $H:\mathbb{R}^{d}\rightarrow\mathbb{R}$:
@@
Figure 8: Potential function whose gradient field is $\operatorname{\tiny\sigma_{\tiny relu}}(\frac{r}{\|x\|}-1)x$, displayed for $x\in\mathbb{R}^{2}$. Repellence force is dynamic: closer to the center (i.e., when a diffusion trajectory is expected to be close to a protected image) it applies stronger gradients, as shown by the arrows, while outside the repellecy radius, it applies no gradient at all, letting the diffusion trajectory continue without any intervention.
@@ Appendix D Implementation Details and Hyperparameters
Input: Batch of latents $\{x_{t}^{(b)}\}_{b=1,\dots,B}$, set of shielded images $\{z_{k}\}_{k=1,\dots,K}$, radius $r$, $\lambda$for $b=1,\dots,B$ do
@@
1        $\triangleright$ Expected diffusion output without repellency
@@
3for $b=1,\dots,B$ do
4       $\vec{\Delta}_{b}\leftarrow 0$for $k=1,\dots,K$ do $\triangleright$ Repel from the shielded set
@@ Appendix E Construction of the Soft-label CC12M Dataset
Appendix F Further Diversity-quality Tradeoffs
In addition to the tradeoff experiments in Section 5.3, Table 4 provides the full combinations of metrics attainable with each method, depending on how one chooses the hyperparameters. This is the raw data underlying Figure 4 and allows the curious reader to compare arbitary tradeoffs.
Table 4: Metrics of all approaches in the tradeoff experiments in Figure 4.
Method Recall Vendi Score Coverage Precision Density FID $FD_{\text{DINOv2}}$ CLIP Score
Base Model 0.237 2.527 0.446 0.558 0.768 9.566 105.967 27.789
Particle Guidance, strength = 1024 0.099 1.987 0.249 0.300 0.326 84.115 705.661 24.470
Particle Guidance, strength = 512 0.230 2.753 0.378 0.443 0.534 23.106 286.093 26.740
Particle Guidance, strength = 256 0.252 2.656 0.429 0.523 0.682 11.934 154.897 27.440
Particle Guidance, strength = 128 0.248 2.591 0.447 0.553 0.754 9.442 109.257 27.704
Particle Guidance, strength = 64 0.245 2.561 0.449 0.559 0.771 9.072 101.796 27.781
Particle Guidance, strength = 32 0.235 2.528 0.445 0.557 0.763 9.724 108.382 27.812
Particle Guidance, strength = 16 0.236 2.529 0.446 0.557 0.764 9.596 107.041 27.813
Interval Guidance, [0.1,0.9] 0.372 2.840 0.455 0.537 0.730 8.385 85.871 27.453
Interval Guidance, [0.2,0.9] 0.419 2.994 0.442 0.514 0.689 8.359 85.094 26.813
Interval Guidance, [0.1,0.8] 0.470 3.174 0.448 0.500 0.663 7.507 76.104 27.215
Interval Guidance, [0.3,0.9] 0.471 3.208 0.421 0.483 0.635 8.406 87.971 25.885
Interval Guidance, [0.2,0.8] 0.518 3.340 0.434 0.478 0.624 7.478 75.250 26.544
Interval Guidance, [0.1,0.7] 0.567 3.576 0.432 0.451 0.577 6.804 72.092 26.784
Interval Guidance, [0.4,0.9] 0.525 3.495 0.395 0.442 0.569 8.623 96.611 24.630
Interval Guidance, [0.3,0.8] 0.571 3.575 0.411 0.446 0.570 7.556 78.887 25.549
Interval Guidance, [0.2,0.7] 0.614 3.770 0.417 0.426 0.536 6.771 72.972 25.979
Interval Guidance, [0.1,0.6] 0.673 4.138 0.396 0.385 0.466 6.885 81.643 26.020
CADS, mixture factor = 0, $\tau_{1}$ = 0.6 0.262 2.598 0.447 0.553 0.753 9.248 105.006 27.746
CADS, mixture factor = 0, $\tau_{1}$ = 0.7 0.253 2.579 0.448 0.555 0.757 9.288 105.549 27.757
CADS, mixture factor = 0, $\tau_{1}$ = 0.8 0.245 2.561 0.449 0.557 0.762 9.356 105.856 27.771
CADS, mixture factor = 0, $\tau_{1}$ = 0.9 0.239 2.545 0.450 0.559 0.767 9.452 106.455 27.790
CADS, mixture factor = 0.001, $\tau_{1}$ = 0.6 0.325 2.816 0.442 0.531 0.696 8.897 105.081 27.534
CADS, mixture factor = 0.001, $\tau_{1}$ = 0.7 0.297 2.734 0.446 0.540 0.719 8.963 104.006 27.617
CADS, mixture factor = 0.001, $\tau_{1}$ = 0.8 0.277 2.660 0.447 0.548 0.739 9.098 103.766 27.697
CADS, mixture factor = 0.001, $\tau_{1}$ = 0.9 0.256 2.588 0.448 0.554 0.755 9.273 105.268 27.754
CADS, mixture factor = 0.002, $\tau_{1}$ = 0.6 0.425 3.208 0.417 0.472 0.584 9.870 129.159 26.920
CADS, mixture factor = 0.002, $\tau_{1}$ = 0.7 0.380 3.028 0.429 0.501 0.637 9.143 114.333 27.242
CADS, mixture factor = 0.002, $\tau_{1}$ = 0.8 0.330 2.837 0.442 0.529 0.692 8.893 105.511 27.506
CADS, mixture factor = 0.002, $\tau_{1}$ = 0.9 0.277 2.660 0.446 0.548 0.739 9.098 103.762 27.696
SPELL, shield radius = 40 0.370 2.998 0.437 0.500 0.631 13.072 140.841 27.397
SPELL, shield radius = 35 0.359 2.935 0.445 0.518 0.665 11.452 120.346 27.556
SPELL, shield radius = 30 0.337 2.856 0.451 0.531 0.695 10.349 106.753 27.655
SPELL, shield radius = 25 0.312 2.774 0.454 0.542 0.723 9.794 100.123 27.739
SPELL, shield radius = 20 0.287 2.691 0.455 0.552 0.746 9.535 98.666 27.781
SPELL, shield radius = 15 0.263 2.616 0.454 0.558 0.762 9.558 100.709 27.811
Appendix G Runtime Analysis and Comparison
The scale of the overhead that SPELL adds is negligible when contrasted with the diffusion generation cost. It amounts to computing (up to) $[B,K]$, distance matrices per time $t$, where both the batchsize $B$ and the size of the protection set $K$ do not exceed hundreds, and adding one single correction vector to the score. Table 5 confirms that the runtime that SPELL adds (as well as the other benchmarked diversity methods) is negligible, here using $B=8$ and intra-batch repellency, hence $K=B-1=7$. This also further confirms that the runtime observed in Section 5.6 is due to the next-neighbor search algorithm, not SPELL’s correction terms.
Table 5: Generation times per image. Neither SPELL nor other diversity inducing methods add considerable runtime. The runtime is dominated by the diffusion backbone. Mean $\pm$ standard deviation across 500 images, run on an NVIDIA V100 GPU.
Model Generation time per image (seconds)
Baseline (Simple Diffusion) 2.93 $\pm$ 0.12
Simple Diffusion + PG 2.96 $\pm$ 0.13
Simple Diffusion + IG 2.93 $\pm$ 0.12
Simple Diffusion + CADS 2.96 $\pm$ 0.12
Simple Diffusion + SPELL 2.94 $\pm$ 0.13
Appendix H Ablation: Repellency Strength Throughout the Generation
@@
Appendix I Image Protection on Large Datasets
@@
Appendix J Examples of Images Generated with Repellency
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Appendix K Ablation: SPELL on Different Prompt Lengths
To verify that SPELL does not only increase the diversity of short prompts, where it is easier to find different images fitting to a prompt, we stratify our analysis by prompt length. Fig. 18 shows that the SPELL increases the diversity, here of the Latent Diffusion model, throughout all prompt length quantiles. Notably, this is on a relatively small repellency radius $r$, which does not decrease the prompt-adherence as measured by the CLIP score.
(a) Recall
(b) CLIP-Score
Figure 18: Diversity and prompt-adherence for short and long prompts. We split the CC12M prompts into 10 categories based on the number of characters. SPELL achieves a consistently higher diversity than the baseline Latent Diffusion model both for short and long prompts, while maintaining the same CLIP-Score as the baseline model. An example for a short prompt is ”Bird on a tree branch” whereas long prompts include ”Head Medusa, creature of Greek mythology. pieces made by hand with goldsmiths and metals such as gold and copper. wears a helmet of green and gold snakes”. Errors bars denote the standard error over 5 seeds.
Appendix L Ablation: Changing the Guidance Weight
@@
Appendix M Ablation: Changing the Repellence Radii
Appendix N Further Diversity Examples
In order to extend Figure 1, we provide further examples of Simple Diffusion without and with SPELL in Figure 27 to Figure 36. The prompts are chosen from MS COCO, which Simple Diffusion was not trained on. As opposed to Figure 1, this features both of SPELL’s capabilities: Intra-batch repellency (every row is a batch of size four), and inter-batch repellency from previous batches, which we treat as the shielded set. The examples affirm qualitatively that SPELL increases the diversity of generated images. Notably, this is without lowering the prompt adherence, which other baselines like IG are prone to, see Table 4 and Figure 4.
We note that some images have copyright overlays, likely learned from the underlying dataset. To investigate this further, we generate 1600 examples for Simple Diffusion without SPELL and 1600 with SPELL. Without SPELL, 62/1600 images have a shutterstock (or similar) overlay, with SPELL it’s 105/1600. To confirm that this is a stable trend, we also generate images with Latent Diffusion (which is trained on the same dataset as Simple Diffusion), where it’s 79/800 without SPELL and 98/800 with SPELL. While the second result could still be a random chance (Chi-Square independence test with Yates’ continuity correction gives p-value = 0.15), the first result is beyond random (p=0.001), and also the effect size is quite measurable (7% vs 4% overlay rate). To improve the understanding of the inner workings, we make two more observations. First, we note that the copyright overlays tend to happen clustered at specific prompts. E.g., one motorcycle prompt has 21/32 images with overlay while most other prompts have 0. So, the distribution is quite skewed and seems to depend on the prompt. Second, we find that the watermarks can serve to push away images from similar ones without the watermark and to pull together images with the same watermark. This informs our best understanding, namely that the copyright overlays serve as a “highway” between modes that allows SPELL to easily explore new modes. SPELL only uses this if the true images / mode distribution of a given prompt actually includes these copyright modes.
(a) Simple Diffusion without SPELL
(b) Simple Diffusion + SPELL
Figure 27: Images generated with Simple Diffusion without and with SPELL for the MS COCO prompt ”A bed and a mirror in a small room.”. Five batches (rows) with each four images, with both intra- and inter-batch repellency, with the same seeds as the runs without SPELL.
(a) Simple Diffusion without SPELL
(b) Simple Diffusion + SPELL
Figure 28: Images generated with Simple Diffusion without and with SPELL for the MS COCO prompt ”Baked pizza with herbs displayed on serving tray at table.”. Five batches (rows) with each four images, with both intra- and inter-batch repellency, with the same seeds as the runs without SPELL.
(a) Simple Diffusion without SPELL
(b) Simple Diffusion + SPELL
Figure 29: Images generated with Simple Diffusion without and with SPELL for the MS COCO prompt ”The clock tower is in the center of the building.”. Five batches (rows) with each four images, with both intra- and inter-batch repellency, with the same seeds as the runs without SPELL.
(a) Simple Diffusion without SPELL
(b) Simple Diffusion + SPELL
Figure 30: Images generated with Simple Diffusion without and with SPELL for the MS COCO prompt ”A bed and desk in a small room.”. Five batches (rows) with each four images, with both intra- and inter-batch repellency, with the same seeds as the runs without SPELL.
(a) Simple Diffusion without SPELL
(b) Simple Diffusion + SPELL
Figure 31: Images generated with Simple Diffusion without and with SPELL for the MS COCO prompt ”A furry, black bear standing in a rocky, weedy, area in the wild.”. Five batches (rows) with each four images, with both intra- and inter-batch repellency, with the same seeds as the runs without SPELL.
(a) Simple Diffusion without SPELL
(b) Simple Diffusion + SPELL
Figure 32: Images generated with Simple Diffusion without and with SPELL for the MS COCO prompt ”A group of seagulls are flying over a wooden dock that is sitting in a lake during the early part of the evening.”. Five batches (rows) with each four images, with both intra- and inter-batch repellency, with the same seeds as the runs without SPELL.
(a) Simple Diffusion without SPELL
(b) Simple Diffusion + SPELL
Figure 33: Images generated with Simple Diffusion without and with SPELL for the MS COCO prompt ”A plane flies over water with two islands nearby.”. Five batches (rows) with each four images, with both intra- and inter-batch repellency, with the same seeds as the runs without SPELL.
(a) Simple Diffusion without SPELL
(b) Simple Diffusion + SPELL
Figure 34: Images generated with Simple Diffusion without and with SPELL for the MS COCO prompt ”A traffic light over a street surrounded by tall buildings.”. Five batches (rows) with each four images, with both intra- and inter-batch repellency, with the same seeds as the runs without SPELL.
(a) Simple Diffusion without SPELL
(b) Simple Diffusion + SPELL
Figure 35: Images generated with Simple Diffusion without and with SPELL for the MS COCO prompt ”a white boat is out on the water”. Five batches (rows) with each four images, with both intra- and inter-batch repellency, with the same seeds as the runs without SPELL.
(a) Simple Diffusion without SPELL
(b) Simple Diffusion + SPELL
Figure 36: Images generated with Simple Diffusion without and with SPELL for the MS COCO prompt ”A table layed out with food such as, salad, steamed peas and carrots, steamed corn, and bread rolls.”. Five batches (rows) with each four images, with both intra- and inter-batch repellency, with the same seeds as the runs without SPELL.